what is the equation in slope-intercept form of the line that has a slope of -3 and contains the point (1, 1)?
Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

what is the equation in slope-intercept form of the line that has a slope of -3 and contains the point (1, 1)?
Plug given information into the "point-slope" form:
y - y1 = m(x - x1)
y - 1 = -3(x - 1)
y - 1 = -3x + 3
y = -3x + 4 (this is what they're looking for)


Answer by stanbon(75887)

what is the equation in slope-intercept form of the line that has a slope of -3 and contains the point (1, 1)?
------
Form: y = mx + b
m= -3 and y = 1 when x = 1
Solve for "b":
1 = -3*1 + b
b = 4
-----
Equation:
y = -3x +4
===============
Cheers,
Stan H.